make " t " the subject of the formula

s=ut+u^2
1
Expert's answer
2012-03-23T14:57:31-0400
To make t the subject means to transform formula into something like t=...
s=ut+u^2
whence ut=s-u^2
therefore t=(s-u^2)/u=s/u-u and that's it